# Heron: Visualizing and Controlling Chemical Reaction Explorations and Networks

### v1

@@ 1 Introduction
For this purpose, we introduce Heron as a graphical user interface (GUI) for our Scine (Software for Chemical Interaction Networks) 9 software project. Heron is open source and free of charge. In addition to the publicly available releases 10, Heron is also part of the AutoRXN workflow 11 on Microsoft Azure and Microsoft Quantum Elements 12, 13. Furthermore, Heron features a graphical integration of the autoCAS 14, 15, 16, 17, 18 software that allows for selecting active orbital spaces in an automated fashion and then automatically carrying out complete active space (CAS) calculations. Moreover, Heron provides interactive real-time quantum mechanics with force feedback which can be applied in research 19, 20, 21, 22, 23 and in educational settings 24, 25. Heron adopts a modular approach where its functionality is encapsulated into different views. For example, the interactive manipulation of a molecule is one module and viewing the results of a rolling automated reaction-mechanism exploration by the software Chemoton 26, 27 is another.
@@
We demonstrate the features of Heron with the example of the Eschenmoser–Claisen rearrangement, which we have recently studied in a kinetically steered first-principles exploration 33. In 1964, Eschenmoser and coworkers discovered that allylic and benzylic alcohols undergo a Claisen rearrangement if heated with N,N-dimethylacetamide dimethyl acetal in xylenes, forming a $\gamma$,$\delta$-unsaturated amide as the product 34. Because this Eschenmoser–Claisen rearrangement is more $(E)$-selective in the formation of the double bond and takes place at lower temperatures ($100-150\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolCelsius}$$C) than other Claisen rearrangements, it has been used in the total synthesis of natural products 35, 36, 37, 38. The key steps of the reaction mechanism of the Eschenmoser–Claisen rearrangement are shown in Figure 1(a). In the first step, the unsaturated alcohol exchanges one of the methoxy groups of N, N-dimethylacetamide dimethyl acetal. The elimination of the second methoxy group and deprotonation of the methyl group follows. Then, the Claisen rearrangement occurs, forming the final product. For an illustrative model system of the reaction, we replaced all side chains RX (X = 1, 2, 3, 4), shown in Figure 1(a), by hydrogen atoms and started the exploration from N,N-dimethylacetamide dimethyl acetal and allyl alcohol. We extracted a reaction mechanism with Heron from the PBE0-D3BJ//GFN2-xTB reaction network of Ref. 33 by considering the concentration fluxes along the reactions based on the visualization presented in section 5 below. The mechanism and the relative free energies encoded in the network are shown in Figure 1(b). In the first two steps of our mechanism, methanol and allyl alcohol are exchanged at the starting acetal 1. Because no proton source is available in toluene, the enamine 2 is formed as an intermediate. The slowest step in the reaction mechanism (with a relative activation free energy $\Delta G^{\ddagger}=182.3\leavevmode\nobreak\ $\mathrm{k}\mathrm{J}\,\mathrm{m}\mathrm{o}\mathrm{l}^{-1}$$) is the dissociation of the second methanol molecule to form the 1,2 and 5,6 unsaturated Claisen rearrangement precursor 4. The [3,3] sigmatropic rearrangement shows a significantly lower barrier than the preceding methanol dissociation of $\Delta G^{\ddagger}=124.8\leavevmode\nobreak\ $\mathrm{k}\mathrm{J}\,\mathrm{m}\mathrm{o}\mathrm{l}^{-1}$$ compared to the starting reactants. We found in Ref. 33 the reaction to be highly exergonic with a reaction free energy of $\Delta G=-125.9\leavevmode\nobreak\ $\mathrm{k}\mathrm{J}\,\mathrm{m}\mathrm{o}\mathrm{l}^{-1}$$.
@@
This work is organized as follows: We discuss structure preparation, running quantum chemical calculations, integration of haptic real-time quantum chemistry, and controlling a rolling reaction network exploration with Heron in sections 2 and 3. We then detail how Heron allows the operator to inspect large network databases, visualize chemical reaction networks, identify paths in large reaction networks, and incorporate information from microkinetic modeling in sections 4 and 5. This is followed by an overview of correlated calculations and the automated selection of active orbital spaces through Heron in section 6 and the construction of interactive hybrid models in section 7. A technical description of its software design is provided in appendix A.
@@
2 Structure Preparation, Manipulation, and Haptic Real-Time Quantum Chemistry
@@
For quantitative studies, structures and paths can be directly transferred to stationary point or minimum energy path search algorithms implemented in ReaDuct 56, 57. In Heron, multiple structures can be stored and subjected to minimization and transition state searches by selection from a pull-down menu. Expert settings for the electronic structure calculations and for the optimization routines can be adjusted in separate pop-up windows. Additionally, more complex optimization workflows (such as structure optimization of reactant and product, then double-ended transition state search, followed by transition state optimization by partitioned rational function optimization 58, and subsequent intrinsic reaction coordinate scans 59) can be constructed by straightforward task chaining. These workflows can also be imported from and exported to human-readable YAML-format files, which can then be run on the command line, allowing a seamless switch between Heron and the command line interface to ReaDuct.
@@
3 Chemical Reaction Networks
3.1 First-principles Exploration
@@
3.2 Activating Filters for Selective Explorations
@@
3.3 Driving an Exploration with a Steering Wheel
@@
One such exploration protocol is provided on Zenodo together with the resulting database 61. It encodes a steered exploration for the Eschenmoser–Claisen rearrangement based on the expected mechanism of SN1 substitution of methanol with the allyl alcohol, second methanol elimination, and final electrocyclic rearrangement. It can reproduce the results of the kinetically-guided exploration in section 5 while lowering the required reaction trials from 19,400 to 12,600 and increasing the number of found nodes corresponding to unique molecular aggregates (compounds and flasks) from 2,220 to 2,411. The steered exploration was set up and executed completely within Heron and required seven Network Expansion Steps.
3.4 Reaction Templates
@@
3.5 Data Management and Statistics
@@ 3.6 Network Visualization
4 Reaction Path Identification
@@
The theory behind encoding kinetic and stoichiometric information in a graph representation of chemical reactions with Pathfinder has been detailed in Ref. 66. In Heron, the Pathfinder interface is accessed from the basic network representation (see section 3.6). A new path analysis window appears, and the operator can enter the IDs of the starting and target compounds in pre-defined fields and adapt the settings of Pathfinder as well. Different types of graphs can be constructed on the fly according to the setting of the operator. One type is the basic graph. In such a graph, all edges from a compound node to a reaction node have the length of 1 and are therefore identical. All edges in the inverse direction, from a reaction node to a compound node, have a length of 0. Hence, in such a graph the length of a path correlates to the number of reactions in this path. Another graph type, the barrier graph, encodes barrier information in edges from a compound node to a reaction node. The higher the barrier of a reaction, the longer the edge from the compound node to this reaction node. The graphs built can be exported and imported. Searching for connections triggers the construction of the underlying graph, basic or barrier, and results in visualizing the first 15 shortest paths between the two nodes of interest, as depicted in Figure 11.
@@
The paths are ordered from top to bottom according to increasing length. The nodes of the paths are represented in the same style as in section 3.6. Each path starts with the starting node (reactant) on the left and ends at the target node (product) on the right (see Figure 11). Between these two end points, reaction and compound nodes occurring along the path are shown (see Figure 11, middle). The compound nodes of incoming reagents required for a specific reaction are shown above the path, to the left of the affected reaction. The compound nodes of outgoing side products of a specific reaction are shown below the path, to the right of the affected reaction.
@@ 5 Microkinetic Modeling
6 Data Refinement by Correlated Calculations with Automated Active Space Selection
@@
7 Interactive Hybrid Models for Nanoscopic Systems
The exploration of chemical structure and reaction space either by a real-time approach described in section 2 or by automated algorithms in Chemoton 26, 27 in section 3.1 becomes computationally challenging for large nanoscale systems such as biomolecules and support-based catalysts. Multiscale models, such as quantum mechanics/molecular mechanics (QM/MM) hybrid models, allow one to hold a “quantum magnifying glass”95 over a local region within a nanoscopic system. In this way, the electronic structure seen in the magnifying glass can be exploited for general reactivity analysis. Application of these models in the context of automated exploratory strategies is hampered by complicated model construction, which we alleviated by automatizing all steps in a single software framework, called Swoose 96, which is a Scine module for multiscale modeling. Heron implements an interface to Swoose.
@@ Acknowledgments
All data newly generated in this work is provided on Zenodo 61.
@@ Appendix A Technical Details
Heron has been primarily written in the Python programming language. Being a widely established, general-purpose programming language, it features both long-term stability and flexibility which are required to maintain Heron over a long time and continuously adapt it to future needs. Furthermore, code development with Python is often faster than with compiled languages such as C++ 102. Moreover, installing a Python package is trivial on most computer systems. However, a well-known drawback of an interpreted language such as Python is that it is rather slow compared to compiled languages such as C++ and Fortran. Fortunately, for most aspects of Heron, Python code runs fast enough on standard contemporary computer hardware. Moreover, Python provides established and reliable mechanisms to offload compute-intensive tasks to very efficient code written, for example, in C++. Many dependencies of Heron take advantage of this; for example, the widely-used Python package NumPy 103 for numerical computations relies on C functions for fast linear algebra operations. Also, Heron contains C++ code for interfacing haptic devices.
Heron relies on several third-party libraries. The most important ones are PySide2 104 (interfacing the well-known Qt software 105 for creating graphical user interfaces), VTK 106 (a toolkit for three-dimensional computer graphics), Matplotlib 107 (which provides functionality for visualizations), SciPy 108 (a library containing many routines for scientific computing), and NumPy 103

### v2

@@ 1 Introduction
For this purpose, we introduce Heron as a graphical user interface (GUI) for our Scine (Software for Chemical Interaction Networks) 9 software project. Heron is open source and free of charge. In addition to the publicly available releases 10, Heron is also part of the AutoRXN workflow 11 on Microsoft Azure and Microsoft Quantum Elements 12, 13. Furthermore, Heron features a graphical integration of the autoCAS 14, 15, 16, 17, 18 software that allows for selecting active orbital spaces in an automated fashion and then automatically carrying out complete active space (CAS) calculations. Moreover, Heron provides interactive real-time quantum mechanics with force feedback which can be applied in research 19, 20, 21, 22, 23 and in educational settings 24, 25. Heron adopts a modular approach where its functionality is encapsulated into different views. For example, the interactive manipulation of a molecule is one module and viewing the results of a rolling automated reaction-mechanism exploration by the software Chemoton 26, 27 is another. In addition to the documentation included in every software release, a video tutorial of the most important functionalities is available at https://youtu.be/HGXkir9iKFo.
@@
This work is organized as follows: We discuss structure preparation, running quantum chemical calculations, integration of haptic real-time quantum chemistry, and controlling a rolling reaction network exploration with Heron in sections 3.1 and 3.2. We then detail how Heron allows the operator to inspect large network databases, visualize chemical reaction networks, identify paths in large reaction networks, and incorporate information from microkinetic modeling in sections 3.3 and 3.4. This is followed by an overview of correlated calculations and the automated selection of active orbital spaces through Heron in section 3.5 and the construction of interactive hybrid models in section 3.6. A technical description of its software design is provided in appendix A.
2 Methods
We demonstrate the features of Heron with the example of the Eschenmoser–Claisen rearrangement, which we have recently studied in a kinetically steered first-principles exploration 33. In 1964, Eschenmoser and coworkers discovered that allylic and benzylic alcohols undergo a Claisen rearrangement if heated with N,N-dimethylacetamide dimethyl acetal in xylenes, forming a $\gamma$,$\delta$-unsaturated amide as the product 34. Because this Eschenmoser–Claisen rearrangement is more $(E)$-selective in the formation of the double bond and takes place at lower temperatures ($100-150\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolCelsius}$$C) than other Claisen rearrangements, it has been used in the total synthesis of natural products 35, 36, 37, 38. The key steps of the reaction mechanism of the Eschenmoser–Claisen rearrangement are shown in Figure 1(a). In the first step, the unsaturated alcohol exchanges one of the methoxy groups of N, N-dimethylacetamide dimethyl acetal. The elimination of the second methoxy group and deprotonation of the methyl group follows. Then, the Claisen rearrangement occurs, forming the final product. For an illustrative model system of the reaction, we replaced all side chains RX (X = 1, 2, 3, 4), shown in Figure 1(a), by hydrogen atoms and started the exploration from N,N-dimethylacetamide dimethyl acetal and allyl alcohol. We extracted a reaction mechanism with Heron from the PBE0-D3BJ//GFN2-xTB reaction network of Ref. 33 by considering the concentration fluxes along the reactions based on the visualization presented in section 3.4 below. The mechanism and the relative free energies encoded in the network are shown in Figure 1(b). In the first two steps of our mechanism, methanol and allyl alcohol are exchanged at the starting acetal 1. Because no proton source is available in toluene, the enamine 2 is formed as an intermediate. The slowest step in the reaction mechanism (with a relative activation free energy $\Delta G^{\ddagger}=182.3\leavevmode\nobreak\ $\mathrm{k}\mathrm{J}\,\mathrm{m}\mathrm{o}\mathrm{l}^{-1}$$) is the dissociation of the second methanol molecule to form the 1,2 and 5,6 unsaturated Claisen rearrangement precursor 4. The [3,3] sigmatropic rearrangement shows a significantly lower barrier than the preceding methanol dissociation of $\Delta G^{\ddagger}=124.8\leavevmode\nobreak\ $\mathrm{k}\mathrm{J}\,\mathrm{m}\mathrm{o}\mathrm{l}^{-1}$$ compared to the starting reactants. We found in Ref. 33 the reaction to be highly exergonic with a reaction free energy of $\Delta G=-125.9\leavevmode\nobreak\ $\mathrm{k}\mathrm{J}\,\mathrm{m}\mathrm{o}\mathrm{l}^{-1}$$.
@@ 2 Methods
3 Results and Discussion
3.1 Structure Preparation, Manipulation, and Haptic Real-Time Quantum Chemistry
@@
For quantitative studies, structures and paths can be directly transferred to stationary point or minimum energy path search algorithms implemented in ReaDuct 56, 57. In Heron, multiple structures can be stored and subjected to minimization and transition state searches by selection from a pull-down menu. Expert settings for the electronic structure calculations and for the optimization routines can be adjusted in separate pop-up windows. Additionally, more complex optimization workflows (such as structure optimization of reactant and product, then double-ended transition state search, followed by transition state optimization by partitioned rational function optimization 58, and subsequent intrinsic reaction coordinate scans 59) can be constructed by straightforward task chaining (see Figure 3). These workflows can also be imported from and exported to human-readable YAML-format files, which can then be run on the command line, allowing a seamless switch between Heron and the command line interface to ReaDuct.
@@
3.2 Chemical Reaction Networks
3.2.1 First-principles Exploration
@@
3.2.2 Activating Filters for Selective Explorations
@@
3.2.3 Driving an Exploration with a Steering Wheel
@@
One such exploration protocol is provided on Zenodo together with the resulting database 61. It encodes a steered exploration for the Eschenmoser–Claisen rearrangement based on the expected mechanism of SN1 substitution of methanol with the allyl alcohol, second methanol elimination, and final electrocyclic rearrangement. It can reproduce the results of the kinetically-guided exploration in section 3.4 while lowering the required reaction trials from 19,400 to 12,600 and increasing the number of found nodes corresponding to unique molecular aggregates (compounds and flasks) from 2,220 to 2,411. The steered exploration was set up and executed completely within Heron and required seven Network Expansion Steps.
3.2.4 Reaction Templates
@@
3.2.5 Data Management and Statistics
@@ 3.2.6 Network Visualization
3.3 Reaction Path Identification
@@
The theory behind encoding kinetic and stoichiometric information in a graph representation of chemical reactions with Pathfinder has been detailed in Ref. 66. In Heron, the Pathfinder interface is accessed from the basic network representation (see section 3.2.6). A new path analysis window appears, and the operator can enter the IDs of the starting and target compounds in pre-defined fields and adapt the settings of Pathfinder as well. Different types of graphs can be constructed on the fly according to the setting of the operator. One type is the basic graph. In such a graph, all edges from a compound node to a reaction node have the length of 1 and are therefore identical. All edges in the inverse direction, from a reaction node to a compound node, have a length of 0. Hence, in such a graph the length of a path correlates to the number of reactions in this path. Another graph type, the barrier graph, encodes barrier information in edges from a compound node to a reaction node. The higher the barrier of a reaction, the longer the edge from the compound node to this reaction node. The graphs built can be exported and imported. Searching for connections triggers the construction of the underlying graph, basic or barrier, and results in visualizing the first 15 shortest paths between the two nodes of interest, as depicted in Figure 11.
@@
The paths are ordered from top to bottom according to increasing length. The nodes of the paths are represented in the same style as in section 3.2.6. Each path starts with the starting node (reactant) on the left and ends at the target node (product) on the right (see Figure 11). Between these two end points, reaction and compound nodes occurring along the path are shown (see Figure 11, middle). The compound nodes of incoming reagents required for a specific reaction are shown above the path, to the left of the affected reaction. The compound nodes of outgoing side products of a specific reaction are shown below the path, to the right of the affected reaction.
@@ 3.4 Microkinetic Modeling
3.5 Data Refinement by Correlated Calculations with Automated Active Space Selection
@@
3.6 Interactive Hybrid Models for Nanoscopic Systems
The exploration of chemical structure and reaction space either by a real-time approach described in section 3.1 or by automated algorithms in Chemoton 26, 27 in section 3.2.1 becomes computationally challenging for large nanoscale systems such as biomolecules and support-based catalysts. Multiscale models, such as quantum mechanics/molecular mechanics (QM/MM) hybrid models, allow one to hold a “quantum magnifying glass”95 over a local region within a nanoscopic system. In this way, the electronic structure seen in the magnifying glass can be exploited for general reactivity analysis. Application of these models in the context of automated exploratory strategies is hampered by complicated model construction, which we alleviated by automatizing all steps in a single software framework, called Swoose 96, which is a Scine module for multiscale modeling. Heron implements an interface to Swoose.
@@ Acknowledgments
The kinetically driven reaction network is provided as part of a larger data set on Zenodo 102. All data newly generated in this work is provided in a separate Zenodo repository 61.
Software Availability Statement
Scine Heron as well as all Scine modules and third-party software mentioned in this manuscript are available open source via the links provided below. All links were accessed on 2024-07-19.
@@ Appendix A Technical Details
Heron has been primarily written in the Python programming language. Being a widely established, general-purpose programming language, it features both long-term stability and flexibility which are required to maintain Heron over a long time and continuously adapt it to future needs. Furthermore, code development with Python is often faster than with compiled languages such as C++ 103. Moreover, installing a Python package is trivial on most computer systems. However, a well-known drawback of an interpreted language such as Python is that it is rather slow compared to compiled languages such as C++ and Fortran. Fortunately, for most aspects of Heron, Python code runs fast enough on standard contemporary computer hardware. Moreover, Python provides established and reliable mechanisms to offload compute-intensive tasks to very efficient code written, for example, in C++. Many dependencies of Heron take advantage of this; for example, the widely-used Python package NumPy 104 for numerical computations relies on C functions for fast linear algebra operations. Also, Heron contains C++ code for interfacing haptic devices.
Heron relies on several third-party libraries. The most important ones are PySide2 105 (interfacing the well-known Qt software 106 for creating graphical user interfaces), VTK 107 (a toolkit for three-dimensional computer graphics), Matplotlib 108 (which provides functionality for visualizations), SciPy 109 (a library containing many routines for scientific computing), and NumPy 104.
@@
Appendix B Exploration Protocol for the Eschenmoser–Claisen Rearrangement
The reaction network is explored step-wise in the KIEA algorithm. First, the algorithm searches for all possible reactions of the educts, and then the concentration flux through the discovered reaction network is simulated by micro-kinetic modeling. KIEA then decides which compounds should be probed for reactions based on the concentration fluxes calculated by the micro-kinetic modeling. This procedure of reaction network exploration, micro-kinetic modeling, and selection of new compounds is continued until no further compounds are trialed for reactions.
Searching for chemical reactions is handled by Chemoton. Chemoton automatically probes compounds for chemical reactions with first-principles calculations, as described in Ref. 26. For these calculations, we modeled the electronic structure with the semi-empirical tight-binding model known as GFN2-xTB52 and the generalized Born and surface area implicit solvation model110, 111 parameterized for GFN2-xTB52 to model toluene as a solvent (referred to as GFN2-xTB(Toluene) in the following sections). Furthermore, we refined the electronic energies for all stationary points on the potential energy surface with density-functional theory (DFT), the exchange–correlation functional known as PBE0112, the def2-TZVP basis set113, the D3 dispersion correction114 with Becke–Johnson damping115, the conductor-like solvation model116 to model toluene, and the quantum chemistry program Turbomole117. The concentration flux through the reaction network is modeled with the Reaction Mechanism Simulator(RMS) micro-kinetic modeling program69, 70, assuming a constant temperature of $150\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolCelsius}$$C, start concentrations for all reactants of $1\leavevmode\nobreak\ $\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{L}^{-1}$$, and propagating for $10\leavevmode\nobreak\ $\mathrm{h}$$. The reaction rate constants $k_{I}^{+/-}$ for the micro-kinetic modeling were calculated according to transition state theory118, 119 as
$k_{I}^{+/-}=\Gamma\frac{k_{B}T}{h}\exp\left[-\frac{\Delta G^{\ddagger+/-}_{I}}{RT}\right]\leavevmode\nobreak\ ,$ (1)
where +/+/-+ / - indicates the direction of the reaction $I$ (forward +, backward -), $\Gamma$ is the transmission coefficient, $k_{B}$ is Boltzmann’s constant, $R$ is the molar gas constant, $T$ is the temperature, $h$ is Planck’s constant, and $\Delta G^{\ddagger+/-}_{I}$ is the free energy of activation of $I$ and its direction at $150\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolCelsius}$$C and $1\leavevmode\nobreak\ $\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{L}^{-1}$$. Furthermore, we assumed $\Gamma=1$ and calculated the free energies from the DFT electronic energies and the commonly used rigid rotor, harmonic oscillator, and particle in a box approximation based on the GFN2-xTB structures. The harmonic vibrational frequencies were calculated with GFN2-xTB(Toluene). We will denote this combination of DFT electronic energies and GFN2-xTB(Toluene) free energy corrections as PBE0-D3BJ(Toluene)//GFN2-xTB(Toluene) in the following.
According to our concentration-flux-steered reaction network exploration algorithm71, we explored unimolecular reactions for all compounds $n$ with an integrated concentration flux $c_{\mathrm{flux}}^{n}$ of more than $1\cdot 10^{-3}\leavevmode\nobreak\ $\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{L}^{-1}$$ and bimolecular reactions between compounds $n$ and $m$ if the product of their maximum concentrations $c_{\mathrm{max}}^{n}c_{\mathrm{max}}^{m}$ exceeded $1\cdot 10^{-2}\leavevmode\nobreak\ $\mathrm{m}\mathrm{o}\mathrm{l}^{2}\,\mathrm{L}^{-2}$$.
[…]
The KIEA algorithm considered no new compounds for the automatic exploration after the sixth micro-kinetic modeling simulation.